Circular No. 7494

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dl IN UGC 1191 M. M. M. Santangelo, Lucca, Italy, reports his discovery of an ap-

parent supernova (R ∼ 18) on five unfiltered CCD frames taken on Sept. 11.05 UT with a 0.51-m reflector in the course of the Monte Agliale Supernovae Search. SN 2000dl was confirmed on nine frames (limiting mag R ∼ 19.5) taken on Sept. 17.89 and 19.90. The new object is located at α = 1^{h}42^{m}05^{s}.77, δ = +7^{o}39′58′′.2 (equinox 2000.0), which is ∼ 4′′ west and 5′′ north of the nucleus of UGC 1191 = MCG +1-5-23 (the northern elliptical galaxy of an interacting pair of galaxies, both of which are near the center of the cluster Abell 240). Nothing appears at this location on the O and E prints of the first Palomar Sky Survey or on various other deep photographic and CCD images of this galaxy.

S. Jha, P. Challis, T. Matheson, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that spectra of SN 2000dl, obtained by P. Berlind on Sept. 21.5 UT with the Whipple Observatory 1.5-m telescope (+ FAST spectrograph), show it to be a type-Ia supernova, 2–3 weeks after maximum light. Cross-correlation of the host-galaxy spectrum with an absorption-line template yields a recession velocity of 18 260 ± 30 km/s.

SUPERNOVA 2000dk IN NGC 382 Jha et al. also report that a spectrum of SN 2000dk (cf. IAUC 7493),

obtained as above on Sept. 20.39 UT, shows it to be a type-Ia supernova at maximum light. The spectrum exhibits a blue continuum, with welldefined Si II, S II, Fe II, Mg II, and Ca II features. Given the NASA/IPAC Extragalactic Database recession velocity of 5228 km/s for the host galaxy, the supernova expansion velocities are 11 500 km/s and 15 500 km/s for Si II (rest 635.5 nm) and Ca II (H and K), respectively. The spectral-feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 2 ± 3 days before maximum light.

XTE J1859+226 P. Zhao, M. Garcia, and J. McClintock, Smithsonian Astrophysical

Observatory, report: “This x-ray nova has declined considerably since July (IAUC 7451, 7463, 7466). The object was not detected in observations made by B. Patten with the F. L. Whipple Observatory’s 1.2-m telescope on Sept. 16–19 UT. The observing conditions were good, and the total exposure times were 30 min for the V band and 45 min for the R band. The magnitude limits are V > 24 and R > 23.”

2000 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green